perimeter of a rectangle is 420 inches.The width is 4 inches less than the length. What is the length?

Answer by mananth(16946)

length = x
width = x-4
..
2*(L+W)=perimeter
2*(x+x-4)=420
2*(2x-4)=420
/2
2x-4 = 210
add +4 to both sides
2x=214
/2
x= 107inches the length
width = 103 inches
...
m.ananth@hotmail.ca